find the base of a triangle if its area is 766 sq. cm. and its altitude is 74 cm.
Found 2 solutions by mananth, shreeaseem007@yahoo.in:
Answer by mananth(16946)

area of triangle = 1/2 *base * height
area =766
height =74
..
766=1/2 * base * 74
766*2/74=base
base 20.7 cm

Answer by shreeaseem007@yahoo.in(3)


A=1/2*b*h
766=1/2*b*74
766/37
this will give you 20.702 answer